Let f : Z → Z be defined by f(a) = 2a 2 − a and g : Z → Z be defined by g(x) = x(2x − 1). Determine whether f is equal to g. Justify your answer.


1
Expert's answer
2021-11-30T12:18:51-0500

We say two functions "f" and "g" are equal if they have the same domain and the same codomain, and if for every "t" in the domain, "f(t)=g(t)."

Let "f : \Z → \Z" be defined by "f(a) = 2a^ 2 \u2212 a."

Domain: "\Z"

Codomain: "\Z"

"\forall t\in\Z:" "f(t)=2t^2-t."


Let "g : \Z → \Z" be defined by "g(x) = x(2x-1)"

Domain: "\Z"

Codomain: "\Z"

"\forall t\in\Z:" "g(t)=t(2t-1)=2t^2-t=f(t)."


Therefore "f" is equal to "g" on "\Z."